OpenStudy (anonymous):

Can someone help me !? Pretty please :D Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (anonymous):

Rectangle 1: length~ 2" height~ 4" width~ 3" volume= 24 (2*4*3= 24) surface area =60 (SA=LA*2B SA=40*2(10) SA=40+20 SA=60) Rectangle 2: (scale *2) length~ 4" height~ 8" width~ 6" volume= 192 (4*8*6=192) surface area= 200 (SA=LA*2B SA=160*2(20) SA=160+40 SA=200)

OpenStudy (anonymous):

what is the problem?

OpenStudy (anonymous):

What is the similarity ratio of these figures along with the ratio of their surface area and volume

OpenStudy (anonymous):

Similarity ratio is the scale

OpenStudy (anonymous):

the scale is 2